# Chain-of-Frames: Advancing Video Understanding in Multimodal LLMs via Frame-Aware Reasoning

### v1

@@ Abstract
Recent work has shown that eliciting Large Language Models (LLMs) to generate reasoning traces in natural language before answering the user’s request can significantly improve their performance across tasks. This approach has been extended to multimodal LLMs, where the models can produce chain-of-thoughts (CoT) about the content of input images and videos. In this work, we propose to obtain video LLMs whose reasoning steps are grounded in, and explicitly refer to, the relevant video frames. For this, we first create CoF-Data, a large dataset of diverse questions, answers, and corresponding frame-grounded reasoning traces about both natural and synthetic videos, spanning various topics and tasks. Then, we fine-tune existing video LLMs on this chain-of-frames (CoF) data. Our approach is simple and self-contained, and, unlike existing approaches for video CoT, does not require auxiliary networks to select or caption relevant frames. We show that our models based on CoF are able to generate chain-of-thoughts that accurately refer to the key frames to answer the given question. This, in turn, leads to improved performance across multiple video understanding benchmarks, for example, surpassing leading video LLMs on Video-MME, MVBench, and VSI-Bench, and notably reducing the hallucination rate. Code available at github.com/SaraGhazanfari/CoF.
@@ 1 Introduction
Large Language Models (LLMs) are able to perform step-by-step reasoning, widely known as chain-of-thoughts (CoT) (Wei et al., 2022; Kojima et al., 2022). This capability has been implicitly integrated into state-of-the-art systems such as OpenAI’s o1/o3 models (OpenAI, 2024) and DeepSeek R1 (Guo et al., 2025), contributing to their remarkable performance and improving the interpretability of their internal functioning. CoT reasoning has been also extended to multimodal LLMs (Hu et al., 2024; Wu and Xie, 2023; Shao et al., 2024): this presents new challenges compared to language-only domains, as the models need to attend to inputs from different modalities, and reason about both their individual content and how they are connected (Awal et al., 2023; Kil et al., 2024; Sun et al., 2025).
@@
Recent work has also begun to explore the integration of CoT into video understanding tasks, where the input to a multimodal LLM consists of a sequence of images (the frames of the video) along with a text prompt. This makes reasoning on videos particularly complex, as the model needs to capture the semantics of the text prompt, understand temporal and causal relationships between frames, and reason about the video in its entirety. Existing approaches rely on complex inference frameworks with architecture modifications (Fei et al., 2024) or auxiliary networks (Han et al., 2024) at evaluation time to integrate reasoning into video LLMs. This makes using these models both more computationally expensive and less general (as they are specialized to some tasks), and deviates from the natural CoT prompting successfully applied to standard LLMs. Another limitation of current methods is that collecting reasoning traces, needed for training, involves complex and costly procedures, which may limit how much the size of training datasets can be scaled in practice. For example, Wang et al. (2024b) iteratively refines the samples via annotations from both LLMs and human experts, Han et al. (2024) leverages several auxiliary models, Fei et al. (2024) uses additional spatial-temporal scene graph data. Finally, video LLMs do not currently provide explicit connections between particular segments of the video and their reasoning, which would better ground the chain-of-thought.
@@
To remedy these limitations, in this work we propose chain-of-frames (CoF), a new frame-aware chain-of-thought reasoning approach for video LLMs that integrates temporal information directly into the CoT structure (see Fig. 1). This enables the model to identify and refer to the most relevant frames while answering questions, in contrast to prior works that treat frame selection and reasoning as separate stages. Chain-of-frames is a simple and natural adaptation of the CoT paradigm in NLP to video understanding that does not require the auxiliary networks or complex inference frameworks of existing methods. Moreover, we propose an efficient data generation pipeline that allows us to collect a large dataset of CoF examples, named CoF-Data. To achieve this, a key element consists of leveraging a synthetic video dataset (Yi et al., 2020) to extract a large and diverse set of reasoning traces at virtually no cost. Then, we fine-tune two recent open-source video LLMs, InternVL2.5-4B (Chen et al., 2024) and InternVL3-8B (Zhu et al., 2025), on our CoF-Data. In an extensive evaluation on five established benchmarks, we show that our CoF models significantly outperform the original ones, with and without naive CoT prompting (Wei et al., 2022) (see Fig. 1). Moreover, they are competitive with or better than state-of-the-art (SOTA) open- and closed-source models (Table 2). Notably, our CoF-InternVL3-8B achieves higher accuracy than the best model reported on Video-MME, MVBench, and VSI-Bench leaderboards, surpassing even proprietary models like GPT-4o and Gemini 1.5 Pro. These results also demonstrate that video LLMs can learn the ability to produce frame-aware reasoning from a limited number of (synthetic) samples and generalize the behavior to unseen tasks. Thus, our work provides a simple and inexpensive approach to adapt the CoT paradigm for video LLMs that improves both performance and interpretability.
(a) A CoF reasoning trace generated by our CoF-4B model.
(b) CoF models vs the baseline models.
Figure 1: (a) Chain-of-frames reasoning generated by our CoF-InternVL2.5-4B model: it includes the key frames to answer the question (from Video-MME). (b) Comparison of accuracy across multiple video understanding benchmarks between baseline models (InternVL2.5-4B and InternVL3-8B) and their CoF-enhanced counterparts: our models consistently outperform the baselines.
@@
Figure 2: CoF-Data. We show examples of our training data with chain-of-frames reasoning, including video, question, answer and reasoning trace. We include samples from CoF-Datareal (real video, top row) and CoF-Datasynth (synthetic video, bottom), created as described in Sec. 3.3.
@@ 2 Related Work
Multimodal LLMs for videos. Multimodal Large Language Models (Chen et al., 2023b; Li et al., 2024b; Ghazanfari et al., 2024) have made substantial progress in integrating visual and textual modalities, enabling them to perform complex reasoning and achieve deep understanding across various data types, including videos (Chen et al., 2023b; Xue et al., 2024; Li et al., 2023a; Zhang et al., 2024b, c; Li et al., 2024a; Bai et al., 2024; Zhu et al., 2025). Among recent advancements, InternVL2.5 (Chen et al., 2024), InternVL3 (Zhu et al., 2025), LLaVA-NeXT-Video (Zhang et al., 2024c), and Qwen2-VL (Wang et al., 2024a) stand out for their strong video understanding capabilities. Both InternVL and LLaVA-NeXT-Video process videos in an image-text interleaved format, aligning sequences of video frames with language to form a unified multimodal stream. In addition to these open-source models, closed-source systems such as GPT-4o (Achiam et al., 2023) and Gemini-1.5 (Gemini Team et al., 2024) have also demonstrated impressive multimodal capabilities, although details of their architecture and training remain proprietary.
Chain-of-Thoughts for videos. Recent research on chain-of-thoughts for video understanding broadly falls into two categories: methods that provide explicit textual reasoning explanations (Wang et al., 2024b), and those that focus on identifying relevant frames to facilitate the process of generating the response (Han et al., 2024; Hu et al., 2025b, a). In the first category, VideoCoT (Wang et al., 2024b) introduces an active annotation tool to generate reasoning explanations, thereby encouraging models to explicitly reason on visual content. In the second category, several approaches emphasize frame selection as a pre-processing step to enhance reasoning efficiency. M-LLM (Hu et al., 2025b) proposes leveraging multimodal LLMs to identify the most relevant frames corresponding to the query. Likewise, Chain-of-Shot (Hu et al., 2025a) introduces a prompting strategy specifically tailored for understanding long-form videos by selecting key frames. VideoEspresso (Han et al., 2024) combines the ideas of core frame selection with fine-grained reasoning annotations, creating a large-scale dataset that supports more efficient and focused video reasoning. Finally, Fei et al. (2024) introduces Video-of-Thought, a complex five-step pipeline to generate spatial-temporal scene graphs to answer multiple-choice questions about videos.
@@ 3.1 Limitations of reasoning on videos
Multimodal LLMs (Zhang et al., 2024c; Chen et al., 2024; Bai et al., 2024) process videos as a sequence of images (frames), which are encoded as image tokens by a vision encoder, then concatenated with the user prompt in natural language, and finally passed to a language model. While the base language models may be trained to produce reasoning traces, these are not specific for reasoning on videos.
To encourage chain-of-thought output in video LLMs, models must be fine-tuned on video-grounded reasoning traces. Wang et al. (2024b) introduces VideoCoT, which used reasoning traces generated by LLMs and refined by human experts to describe video events. However, the traces lack explicit temporal grounding, that is, individual reasoning steps are not clearly aligned with the corresponding video frames. Additionally, the reliance on human and LLM annotations makes data generation expensive, limiting the dataset to only 11k samples. A different strategy is proposed by Han et al. (2024), who use multiple auxiliary models to generate CoT data. Lightweight multimodal LLMs select the core frames, GPT-4o (Hurst et al., 2024) identifies key elements in each core frame and their relevance to the query, GroundingDINO (Liu et al., 2024) provides spatial annotations (e.g., bounding boxes), and the BGE-M3 retriever (Chen et al., 2023a) produces the temporal annotations. While effective during training, this approach cannot be deployed at inference time given its high complexity and cost (only the frame selection step, which depends on auxiliary LLMs, is kept). Alternatively, Video-of-Thought (Fei et al., 2024) introduces a complex five-step inference pipeline, which includes generating the spatial-temporal scene graphs for key frames and is specialized for multiple-choice questions.
In summary, existing approaches to video reasoning with LLMs face three main limitations: (i) expensive training data generation, (ii) complex inference, possibly involving auxiliary models, and (iii) lack of explicit temporal grounding in the reasoning process.
@@ 3.2 Chain-of-Frames
To address these limitations, we propose Chain-of-Frames (CoF), a simple yet effective approach introducing temporal grounding into the reasoning process. CoF consists of reasoning traces with explicit references to frames relevant to answering the given query. Concretely, we use the position of the frame in the video (e.g., “Frame 1”, “Frame 2”, …) as an identifier. Unlike timestamps, this representation is agnostic to video duration and sample frequency, making it more consistent across diverse video data and potentially easier to learn. Examples are shown in Fig. 1 and Fig. 2, with additional illustrations provided in App. C. This approach has several benefits
@@
    Data availability: as we show in Sec. 3.3, training data can be efficiently generated from existing annotated real videos. Additionally, CoF traces can be created from synthetic videos at virtually no cost. This low-cost data generation contrasts with the more complex pipelines of Wang et al. (2024b); Han et al. (2024).
@@
    Simplicity: the reasoning traces are completely in natural language (without the need for complex format, bounding boxes or scene graphs as for Fei et al. (2024); Han et al. (2024)). This approach is a natural extension of standard CoT method used in NLP tasks and does not require specific inference pipelines.
@@
    Temporal grounding: by explicitly referencing the frames relevant to the task, CoF increases the connection between video and reasoning. Thus, our method can be seen as a temporally grounded extension of the typical CoT of Wang et al. (2024b).
@@
    Interpretability: the reasoning traces generated at inference time include frame references (Fig. 1) and provide direct insights into how the LLM’s decisions are obtained.
In the next section, we detail how we generate chain-of-frames from video datasets. These traces constitute the training data for fine-tuning video LLMs capable of producing CoF-based reasoning.
Figure 3: Overview of our two-step pipeline for generating CoF-Data. Step 1 adjusts the frame IDs while preserving frame-caption alignment. Step 2 utilizes raw annotations to generate CoF triplets (question, frame-aware reasoning trace, answer). For this, we leverage Llama3.1-8B a manual template with the synthetic videos from Clevrer.
3.3 Chain-of-Frames training data collection (CoF-Data)
We construct chain-of-frames traces from both real and synthetic videos. For real videos, we use the training split of the VideoEspresso dataset (Han et al., 2024), which features videos from a wide range of sources and includes descriptions of key frames for each. These diverse frame-level annotations provide a starting point for extracting CoF traces. To complement this, we use the training split of the Clevrer dataset (Yi et al., 2020), which contains synthetic videos of simple 3D objects interacting within a controlled environment, along with rich annotations. Synthetic data offers two key advantages: it overcomes the cost of generating training samples from real videos and mitigates the scarcity of sufficiently annotated real videos. We next describe the main steps of our data generation pipeline (illustrated in Fig. 3), and introduce the resulting dataset of chain-of-frames, named CoF-Data. Additional details and examples are provided in App. A.
@@ 3.3 Chain-of-Frames training data collection (CoF-Data)
CoF from real videos (CoF-Datareal). VideoEspresso provides captions for key frames. After aligning frame IDs, we obtain data in the format shown in Fig. 3 (raw annotations). From these annotations, we generate triplets of questions, answers, and reasoning traces with frame references by prompting an LLM using the raw annotations as input. For this, we use Llama3.1-8B-Instruct (Meta AI, 2024) (the full prompt is provided in App. A). This process yields multiple questions per video, often covering diverse parts of the video and referencing different sets of frames. An example from CoF-Datareal is illustrated in Fig. 3, and a complete training sample including the video is in Fig. 2.
CoF from synthetic videos (CoF-Datasynth). In each frame of Clevrer, every object is annotated with both fixed properties (shape, material, color) and situational attributes (e.g., velocity, location), see Fig. 3. We use these rich attributes to generate three categories of quantitative questions (object count, appearance order, and relative distance) which complement the semantic questions obtained for real video. Notably, we can generate both questions, answers, and chain-of-frames using fixed manual templates (see App. A), since all the necessary information can be directly deduced from the object-specific raw annotations. This eliminates the need for using an LLM, significantly reducing generation cost and enabling easy scaling of the dataset size. An example of an appearance order question is shown in Fig. 2 (examples from other categories are shown in App. C).
Final dataset (CoF-Data). From the generated chain-of-frames, we filter out samples where frames are referred in the question (as this does not happen at test time). Moreover, we reduce the number of samples with no frame references in the reasoning trace to give higher weight to more complex examples of reasoning. We nevertheless keep a non-negligible fraction of samples with no frame references since there might be, in the evaluation benchmarks, questions which do not require CoF-like reasoning, and we do not want to force the model to generate it when unnecessary. This yields a total of 164,186 samples, comprising 103,683 samples from the CoF-Datareal dataset, which is based on real-world videos, and 60,503 samples from the CoF-Datasynth dataset of synthetic videos. Fig. 4 shows the distribution of how many frames are referenced in the reasoning traces, both for the final dataset and the individual splits. The CoF-Datasynth exhibits a more balanced distribution compared to the automatically generated CoF-Datareal: this highlights that using synthetic videos allows us to better control various aspects of the data.
Figure 5: Video encoding format. InternVL models add a textual identifier before each frame, which is well-suited for our chain-of-frames reasoning. The illustration is taken from Chen et al. (2024).

### v2

@@ Abstract
Recent work has shown that eliciting Large Language Models (LLMs) to generate reasoning traces in natural language before answering the user’s request can significantly improve their performance across tasks. This approach has been extended to multimodal LLMs, where the models can produce chains-of-thoughts (CoT) about the content of input images and videos. For video inputs, prior works use complex multi-step pipelines that extract and include relevant frames from videos in the CoT, or produce simpler single-stage reasoning traces at the expense of poor temporal grounding. Here, we propose the first video LLMs with single-stage reasoning that includes explicit references to relevant frames, thereby reducing temporal inconsistencies in the reasoning process. Our approach is simple, unified, and self-contained, employing a single-stage inference to handle complex video understanding tasks without relying on auxiliary modules for frame selection or caption generation. For this, we first create CoF-Data, a large dataset of diverse questions, answers, and corresponding frame-grounded reasoning traces from both natural and synthetic videos, spanning various topics and tasks. Our models, obtained fine-tuning video LLMs on this chain-of-frames (CoF) data, generate reasoning traces that accurately identify key frames to answer given questions. In turn, this consistently improves performance across multiple video understanding benchmarks. Surprisingly, we find that synthetic data alone, despite being out-of-distribution with respect to these real-world benchmarks, provides a significant boost in model accuracy. Code available at GitHub.
(a) A Chain-of-Frames reasoning trace generated by our CoF-4B model.
(b) CoF models vs the baseline models.
Figure 1: (a) Chain-of-Frames (CoF) reasoning generated by our CoF-InternVL2.5-4B model, referencing the key frames to answer the question (from Video-MME). (b) Comparison of accuracy across multiple video understanding benchmarks between baseline models (InternVL2.5-4B and InternVL3-8B) and their CoF-enhanced counterparts: our models outperform the baselines in nearly every case.
@@ 1 Introduction
Large Language Models (LLMs) are able to perform step-by-step reasoning, widely known as chain-of-thoughts (CoT) [34, 21]. This capability has been implicitly integrated into state-of-the-art systems such as OpenAI’s o1/o3 models [28] and DeepSeek R1 [13], contributing to their remarkable performance and improving the interpretability of their internal functioning. CoT reasoning has been also extended to multimodal LLMs [18, 35, 29]: this presents new challenges compared to language-only domains, as the models need to attend to inputs from different modalities, and reason about both their individual content and how they are connected [3, 20, 30].
@@
Recent work has also begun to explore the integration of CoT into video understanding tasks, where the input to a multimodal LLM consists of a sequence of images (the frames of the video) along with a text prompt. This makes reasoning on videos particularly complex, as the model needs to capture the semantics of the text prompt, understand temporal and causal relationships between frames, and reason about the video in its entirety. Existing approaches rely on complex inference frameworks with architecture modifications [9] or auxiliary networks [14] at evaluation time to extract and integrate key frames from the video sequence into the reasoning trace. This makes using these models both more computationally expensive and less general (as they are specialized to some tasks), and deviates from the natural CoT prompting successfully applied to standard LLMs. Another limitation of current methods is that collecting reasoning traces, needed for training, involves complex and costly procedures, which may limit how much the size of training datasets can be scaled in practice. For example, Wang et al. [32] iteratively refine reasoning traces via annotations from LLMs and human experts, Han et al. [14] leverage several auxiliary models, and Fei et al. [9] use additional spatial-temporal scene graph data. Finally, video LLMs do not currently provide explicit connections between particular segments of the video and their reasoning, which would better ground the chain-of-thought.
@@
To remedy these limitations, in this work we propose Chain-of-Frames (CoF), a new frame-aware chain-of-thought reasoning approach for video LLMs that integrates temporal information directly into a single-shot text-only CoT structure (see Fig. 1a). Chain-of-frames is a simple and natural adaptation of the CoT paradigm in NLP to video understanding that does not require the auxiliary networks or complex inference frameworks of existing methods. Moreover, we propose an efficient data generation pipeline that allows us to collect a large dataset of CoF examples, named CoF-Data. To achieve this, a key element consists of leveraging a synthetic video dataset [39] to extract a large and diverse set of temporally grounded reasoning traces at virtually no cost. Then, we fine-tune three recent open-source video LLMs, InternVL2.5-4B [6], InternVL3-8B [43] and Phi-3.5-Vision [1] on our CoF-Data. In an extensive evaluation on five established benchmarks, we show that our CoF models significantly outperform the original ones with and without naive CoT prompting [34] (see Fig. 1b). Moreover, they are competitive with or better than state-of-the-art (SOTA) open- and closed-source models (Tab. 1). Finally, further analysis reveals that synthetic data alone—despite being out-of-distribution relative to real-world benchmarks—provides a substantial improvement in model accuracy. This indicates that the model effectively learns the underlying task from the synthetic data and successfully transfers this knowledge to real-world scenarios.
@@
Figure 2: CoF-Data. We show examples of our training data with chain-of-frames reasoning, including video, question, reasoning trace and answer. We include samples from CoF-Datareal (real video, top row) and CoF-Datasynth (synthetic video, bottom), created as described in Sec. $3.$
@@ 2 Related Work
Multimodal LLMs for videos. Multimodal Large Language Models [7, 25, 12] have made substantial progress in integrating visual and textual modalities, enabling them to perform complex reasoning and achieve deep understanding across various data types, including videos [7, 37, 22, 41, 42, 23, 4, 43]. Among recent advancements, InternVL2.5 [6], InternVL3 [43], LLaVA-NeXT-Video [42], and Qwen2-VL [31] stand out for their strong video understanding capabilities. Both InternVL and LLaVA-NeXT-Video process videos in an image-text interleaved format, aligning sequences of video frames with language to form a unified multimodal stream. In addition to these open-source models, closed-source systems such as GPT-4o [2] and Gemini-1.5 [11] have also demonstrated impressive multimodal capabilities, although details of their architecture and training remain proprietary.
Chain-of-Thoughts for videos. Recent research on chain-of-thought (CoT) reasoning for video understanding can be broadly categorized into two paradigms: (1) single-stage methods, which directly generate textual reasoning explanations followed by an answer in a single step [32], and (2) multi-stage methods, which employ auxiliary models to extract meta-information in earlier stages and then use this information to generate the answer in the final step [14, 17, 16]. In the first category, VideoCoT [32] introduces an active annotation framework to produce reasoning explanations from video descriptions, thereby guiding the model to generate reasoning traces followed by the answer. In the second category, most approaches [33, 17, 16, 14, 5] emphasize key frame selection as an initial stage to enhance downstream answer generation. Building on this idea, VideoEspresso [14] further integrates additional steps, combining the selected core frames with the question and feeding them into a reasoning model that performs chain-of-thought (CoT) reasoning, with the final answer produced in a subsequent stage. Finally, Video-of-Thought [9] introduces a more complex five-stage pipeline that constructs spatio-temporal scene graphs to reason over videos and answer multiple-choice questions.
@@ 3.1 Limitations of reasoning on videos
Recent work has begun to explore the integration of CoT into video understanding tasks. While the base language models may be trained to produce reasoning traces, these are not specific for reasoning on videos. To encourage chain-of-thought output in video LLMs, models must be fine-tuned on video-grounded reasoning traces. VideoCoT [32] addresses this by generating CoT training data through a pipeline that leverages video descriptions as raw annotations and uses LLMs to produce corresponding reasoning traces. However, because the descriptions are not temporally aligned at the frame level, the resulting reasoning traces can introduce temporal inconsistencies [24]. To mitigate such inaccuracies, human annotators are introduced into the loop to refine the LLM-generated reasoning traces. As a result, the reliance on human and LLM annotations makes data generation expensive, limiting the dataset to only 11k samples. Moreover, the traces lack explicit temporal grounding; that is, individual reasoning steps are not clearly aligned with the corresponding video frames.
A different strategy is to leverage multiple auxiliary models to generate the meta-information, including finding the key frames from the video and explicitly grounding the model’s answer to the key frames [14, 17, 9]. While these multi-stage pipelines improve video reasoning, their computational overhead during inference poses a major bottleneck. Furthermore, because only a subset of the frames is passed to the video LLM, the video’s full temporal context is lost.
In summary, current video reasoning methods face two major limitations: (1) they depend on expensive and often inaccurate training data generation while lacking explicit temporal grounding in the reasoning process, or (2) involve complex inference procedures that rely on auxiliary models and compromise contextual understanding by processing only a subset of the frames to generate the final answer.
Figure 3: Overview of our two-step pipeline for generating CoF-Data. Step 1 adjusts the frame IDs while preserving frame-caption alignment. Step 2 utilizes raw annotations to generate CoF triplets (question, frame-aware reasoning trace, answer), using Llama-3.1-8B-Instruct for real-video data and manual templates for synthetic data.
@@ 3.2 Chain-of-Frames
To address these limitations, we propose Chain-of-Frames (CoF), a simple yet effective approach introducing temporal grounding into the reasoning process. CoF consists of reasoning traces with explicit references to frames relevant to answering the given query, within a single-stage inference. Concretely, we use the position of the frame in the video (e.g., “Frame 1”, “Frame 2”, …) as an identifier. Unlike timestamps, this representation is agnostic to video duration and sample frequency, making it more consistent across diverse video data and potentially easier to learn. Examples are shown in Fig. 1 and Fig. 2, with additional illustrations provided in App. C. This approach has several benefits:
@@
    Data quality: as we show in Sec. 3.3, training data can be efficiently generated from existing annotated real videos, where the video descriptions are temporally aligned at the frame level, resulting in temporally accurate reasoning traces. Moreover, CoF traces can be generated from synthetic videos with high precision by combining their rich annotations with manually designed templates, at virtually no additional cost. This low-cost and scalable data generation process stands in contrast to the more complex pipelines [32, 14].
@@
    Simplicity: our reasoning traces are expressed entirely in natural language—without requiring auxiliary modules to generate complex data formats, bounding boxes, or scene graphs as in Fei et al. [9], Han et al. [14]. This makes our approach a natural extension of standard Chain-of-Thought (CoT) reasoning used in NLP tasks, eliminating the need for multi-stage inference pipelines and reducing overall latency.
@@
    Explicit temporal grounding: temporal grounding is achieved by explicitly referencing the task-relevant frames within the reasoning traces, rather than by retrieving and processing key frames through a multi-stage inference process. This design strengthens the connection between visual content and reasoning, making CoF a temporally grounded extension of the typical CoT approach proposed by Wang et al. [32].
@@
    Interpretability: The reasoning traces generated at inference time include frame references (Fig. 1), providing direct insights into how the LLM’s decisions are obtained.
Figure 4: Video encoding format. InternVL models add a textual identifier before each frame, which is well-suited for our chain-of-frames reasoning. The illustration is taken from Chen et al. [6].
3.3 Chain-of-Frames training data collection
We construct chain-of-frames traces from both real and synthetic videos. For real videos, we use the training split of the VideoEspresso dataset [14], which features videos from a wide range of sources and includes descriptions of key frames for each. These diverse frame-level annotations provide a starting point for extracting CoF traces. To complement this, we use the training split of the Clevrer dataset [39], which contains synthetic videos of simple 3D objects interacting within a controlled environment, along with rich annotations. Synthetic data offers two key advantages: it overcomes the cost of generating training samples from real videos and provides highly accurate temporally grounded reasoning trace derived from the rich annotations with manual templates. We next describe the main steps of our data generation pipeline, illustrated in Fig. 3. Additional details and examples are provided in App. A.
@@ 3.3 Chain-of-Frames training data collection
CoF from real videos (CoF-Datareal). VideoEspresso provides captions for key frames. After aligning frame IDs, we obtain data in the format shown in Fig. 3 (raw annotations). From these temporally aligned annotations, we generate triplets of questions, answers, and reasoning traces with frame references by prompting an LLM using the raw annotations as input. For this, we use Llama3.1-8B-Instruct [27] (the full prompt is provided in App. A). This process yields multiple questions per video, often covering diverse parts of the video and referencing different subsets of frames. An example from CoF-Datareal is in Fig. 3, and a complete training sample including the video is in Fig. 2.
CoF from synthetic videos (CoF-Datasynth). In each frame of Clevrer, every object is annotated with both fixed properties (shape, material, color) and situational attributes (e.g., velocity, location), see Fig. 3. We use these rich attributes to generate three categories of quantitative questions (object count, appearance order, and relative distance) which complement the semantic questions obtained for real video. Notably, we can generate both questions, answers, and chain-of-frames using fixed manual templates (see App. A), since all the necessary information can be directly deduced from the object-specific raw annotations. This eliminates the need for using an LLM, significantly reducing generation cost and enabling easy scaling of the dataset size. An example of an appearance order question is shown in Fig. 2 (for other categories see App. C).
Final dataset (CoF-Data). From the generated chain-of-frames, we filter out samples where frames are referred in the question (as this does not happen at test time). Moreover, we reduce the number of samples with no frame references in the reasoning trace to give higher weight to more complex examples of reasoning. We nevertheless keep a non-negligible fraction of samples with no frame references since there might be, in the evaluation benchmarks, questions which do not require CoF-like reasoning, and we do not want to force the model to generate it when unnecessary. The resulting dataset comprises 164,186 samples. Fig. 5 illustrates the distribution of the number of frames referenced per sample.
Figure 5: Distribution of frame references in CoF-Data.